# Boosting Open Set Recognition Performance through Modulated Representation Learning

### v1

@@ 3.2.1 Generalized Cosine Scheduling
Instead of a constant value, we propose to schedule the temperature from a range $[\tau^{-},\tau^{+}]$ as the training progresses using a generalized cosine scheduling, which is defined as
$\text{GCosSch}(e;\tau^{+},\tau^{-},P,k)=\left\{\begin{array}[]{ll}\tau^{-}+\frac{1}{2}(\tau^{+}-\tau^{-})(1+\text{cos}(\frac{2\pi e}{P}-k\pi)),&\mbox{if }e\leq E-\frac{kT}{2}\\ \tau^{+},&\mbox{elsewhere }\end{array}\right.$ (4)
@@
where value $k\pi$ represents the delay of the cosine wave with respect to the starting epoch, $P$ is the period of the wave (Figure 2) and $k$ can be from $[0,1]$. $k=0$ reduces (4) to the regular cosine TS (CosSch), proposed by Kukleva et al. (2023) for the self-supervised tasks on the long-tailed datasets. With CosSch, the model starts training with a higher temperature $\tau^{+}$ and goes to a lower temperature $\tau^{-}$.
$\text{CosSch}(e;\tau^{+},\tau^{-},P)=\text{GCosSch}(e;\tau^{+},\tau^{-},P,k=0)$ (5)
@@
However, we find that rather than using (5), it is beneficial for the task switching if we start with a lower temperature $\tau^{-}$ and move towards a higher value $\tau^{+}$. Starting with a lower temperature, the model provides priority to fewer neighbors, learning the coarse structure of representation space, resulting in a sharper decision boundary. The open set samples remain distant from any cluster of the known classes (due to heavy penalization of slight dissimilarity). As the temperature increases, the model prioritizes more neighbors and gradually pulls the positive samples to its own cluster. This makes within-class representations more compact and the decision boundary smoother while the core separation learned earlier is maintained. The open set samples are not pulled as tightly because the features are unknown to the model, maintaining the separation. This leads to a richer and potentially more generalized representation space for both open set and closed set performance. Figure 1 illustrates the concept with UMAP projection of representation spaces for different temperature schedules.
@@
For $k=1$ in (4), the temperature starts with a lower value and goes to a higher one, looking like a negative cosine wave, hence the name negative cosine scheduling (NegCosSch).
$\text{NegCosSch}(e;\tau^{+},\tau^{-},P)=\text{GCosSch}(e;\tau^{+},\tau^{-},P,k=1)$ (6)
@@
Through experimental results, we find that NegCosSch is better for OSR than CosSch or any other value of $k$ in GCosSch. The model settles down better if the final few epochs maintain a higher temperature rather than follow the wave (epochs 500-600 in Figure 2). Although initially aimed for OSR, the NegCosSch also improves performance for closed set classification while incurring no additional computational burden as it only includes an epoch-dependent temperature in a loss function.
3.3 Inference
@@
For CE loss, we use the model as is for inference. However for SupCon loss, we remove the projection layer and a linear classifier is trained for evaluation. We use the maximum logit based scoring rule for OSR score.
@@ 4 Results and Discussion
Benchmarks: We test our method on four regular OSR benchmarks: CIFAR10, CIFAR+10, CIFAR+50, and TinyImageNet. Moreover, we test our method on the SSBs defined on three fine-grained datasets: the Caltech-UCSD-Birds (CUB) (Wah et al. (2011)), FGVC-Aircraft (Krause et al. (2013)) and Stanford Cars (SCars) (Maji et al. (2013)). For these benchmarks, the open set classes are divided into ‘Easy’ and ‘Hard’ splits by computing the semantic similarity (based on the labeled visual attributes) of each pair of classes, the details of which can be found in Vaze et al. (2022). The different difficulty level along with more training classes make these datasets harder OSR benchmarks than the regular ones.
(a) CIFAR10
(b) CIFAR+10
(c) CIFAR+50
(d) TinyImageNet
Figure 3: Open Set AUROC of different temperature scheduling for the SupCon loss on the regular OSR benchmarks.
@@
Training Details: We mostly follow the experiment settings and design choices from Vaze et al. (2022). For the regular benchmarks, we use a VGG32-like model and for the SSB datasets, we use a ResNet50 model pretrained on the places365 dataset [^2] In spite of our efforts, we could not find the same pretrained model mentioned in Vaze et al. (2022) online. Therefore, we use the pretrained model from Zhou et al. (2017) trained on places365, which is completely unrelated to the SSBs..
For experiments, we choose temperatures from $\mathcal{T}_{SupCon}=\{0.025,0.05,0.1,0.2,0.3,0.4,0.5\}$ for the SupCon loss and from $\mathcal{T}_{CE}=\{0.5,1.0,2.0\}$ for the CE loss. For label smoothing (LS) in SupCon loss, we implement the following function instead of (2):
with $N_{i}(\alpha)=\sum_{k\in I\setminus\{i\}}[(1-\alpha)\mathbf{1}_{y_{i}=y_{k}}+\frac{\alpha}{C-1}\mathbf{1}_{y_{i}\neq y_{k}}]$, where $\alpha$ is the smoothing coefficient and $\mathbf{1}$ is the indicator function. The training details are provided in the Appendix. All the experiments are performed on a single NVIDIA RTX2080Ti GPU requiring from 2 to 32 hours to train a single model depending on the model and the dataset.
@@
Metrics: We report the closed set performance as a $C$-class classification using accuracy (%), the open set performance as known-unknown detection using AUROC (%), and the open set classification rate (OSCR %). OSCR measures the trade-off between the open set and the closed set performance (Dhamija et al. (2018)). Now, we discuss the results.
4.1 NegCosSch yields better OSR performance (AUROC) than both CosSch and constant temperatures
Figure 3 presents the OSR performance of our proposed TS scheme along with the regular cosine scheduling and constant temperatures in SupCon loss on the regular benchmarks. We vary $(\tau^{+},\tau^{-})$ from $\mathcal{T}^{2}_{SupCon}$ and compare $\text{NegCosSch}(\tau^{+},\tau^{-},P)$ with $\text{CosSch}(\tau^{+},\tau^{-},P)$, $\text{Const}(\tau=\text{nearest}(\frac{1}{2}[\tau^{+}+\tau^{-}]))$ [^3] By ‘nearest’, we mean a nearest temperature is chosen from $\mathcal{T}_{SupCon}$. and $\text{Const}(\tau=\tau^{-})$. The objective of the quadruplet-wise comparisons is to determine if our proposed scheduling method outperforms a regular cosine scheduling, a constant temperature set to the midpoint of $(\tau^{+},\tau^{-})$, or set to $\tau^{-}$. We observe that for the CIFAR10, CIFAR+10 and TinyImageNet benchmarks, our proposed scheduling yields a better open set AUROC than CosSch and constant temperatures for most of the quadruplet comparisons. We find the improvement or degradation to be insignificant for the CIFAR+50 benchmark with the highest AUROC found for CosSch($0.3,0.1$).
Specifically, for any good value of constant temperature $\tau$ (which can be chosen from hyperparameter tuning) in the SupCon loss, it is better to use NegCosSch by setting $\tau^{+}=\tau+\Delta,\tau^{-}=\tau-\Delta$ (or alternatively, $\tau^{+}=\tau+\Delta,\tau^{-}=\tau$) with the increment Δ0.1 or $\Delta\approx 0.1\text{~{}or~{}}0.2$. For example, in TinyImageNet, we find from hyperparameter tuning that Const($\tau=0.2$) provides a good OSR performance. Therefore, NegCosSch($\tau^{+}=0.4,\tau^{-}=0.1$) or NegCosSch($\tau^{+}=0.3,\tau^{-}=0.2$) are better choices than Const($\tau=0.2$), CosSch($\tau^{+}=0.4,\tau^{-}=0.1$) and CosSch($\tau^{+}=0.3,\tau^{-}=0.2$). Utilizing this, we reduce the search space of $(\tau^{+},\tau^{-})$ on the SSB experiments, where we find higher performance boost discussed in Section 4.3. The closed set performance and AUROC with other values of $k$ in GCosSch are listed in the Appendix.
4.2 Improving OSR Methods with Our Proposed Scheme
Here to evaluate our scheme, we incorporate NegCosSch into three OSR methods: the CE baseline by Vaze et al. (2022), the SupCon baseline by Khosla et al. (2020) and the ConOSR by Xu et al. (2023) on the regular benchmarks. We improve the baseline scores of ConOSR for TinyImageNet with better design choices. The results shown in Table 1 indicate that adding our TS on top of these methods results in a similar or higher open set AUROC for most cases. We suspect that the benefits of our TS reduce when the number of training classes is small. For example, there are only 6 training classes in CIFAR10 and 4 training classes in the CIFAR+10 and CIFAR+50 benchmarks. For TinyImageNet with 20 training classes, our scheme shows significant improvements for all methods. We now consider the SSBs, where they have a larger number of training classes.
Table 1: Open set performance (AUROC %) comparison of different OSR methods with and without our proposed scheme.
Methods CIFAR10 CIFAR+10 CIFAR+50 TinyImageNet
CE (Baseline) (Vaze et al. (2022)) 93.6 97.9 96.5 83
CE + NegCosSch (ours) 93.76 97.88 96.32 83.47
SupCon (Baseline) 94.04 98.25 97.03 82.86
SupCon + NegCosSch (ours) 94.10 98.28 97.03 83.09
ConOSR (Xu et al. (2023)) 94.55 98.22 97.04 83.40
ConOSR + NegCosSch (ours) 94.52 98.29 97.01 83.63
Table 2: Performance on the SSBs. Open set results are shown on ‘Easy / Hard’ splits. Here, NC-TS means our negative cosine temperature scheduling and w/o LS means without label smoothing. We highlight the cases where our scheduling outperforms the corresponding baseline.
CUB FGVC-Aircraft SCars
Methods Acc. AUROC OSCR Acc. AUROC OSCR Acc. AUROC OSCR
ARPL (Chen et al. (2021)) 85.9 83.5 / 75.5 76.0 / 69.6 91.5 87.0 / 77.7 83.3 / 74.9 96.9 94.8 / 83.6 92.8 / 82.3
ARPL + NC-TS(ours) 85.97 87.46 / 79.5 79.38 / 73.5 90.58 91.06 / 81.40 85.76 / 77.46 96.97 95.82 / 86.24 93.83 / 84.95
CE (w/o LS) 84.60 83.40 / 74.68 70.59 / 63.29 90.29 91.33 / 81.75 82.38 / 73.95 96.66 93.20 / 84.32 90.17 / 81.70
CE (w/o LS) + NC-TS(ours) 86.20 86.96 / 77.67 74.91 / 67.01 90.95 91.82 / 83.33 83.57 / 76.1 97.49 94.59 / 86.22 92.25 / 84.14
CE + LS (Vaze et al. (2022)) 85.71 85.15 / 77.19 72.93 / 66.19 90.65 86.99 / 78.85 78.94 / 71.72 96.96 94.33 / 84.00 91.54 / 81.62
CE + LS + NC-TS(ours) 86.34 86.86 / 77.86 74.92 / 67.29 91.19 87.48 / 79.86 79.84 / 73.07 97.11 94.50 / 84.02 91.82 / 81.75
SupCon (w/o LS) 83.22 86.50 / 73.64 71.93 / 61.34 89.63 85.33 / 81.03 76.74 / 73.05 96.88 93.00 / 82.77 90.19 / 80.32
SupCon(w/oLS) + NC-TS(ours) 83.95 87.35 / 74.60 73.38 / 62.84 88.49 86.56 / 81.23 76.84 / 72.36 96.66 93.16 / 82.82 90.14 / 80.19
SupCon + LS 83.88 86.30 / 73.17 72.28 / 61.38 88.73 86.94 / 81.40 77.42 / 72.68 96.58 93.22 / 82.88 90.10 / 80.16
SupCon+LS+ NC-TS(ours) 83.81 86.49 / 74.40 72.38 / 62.34 89.57 85.36 / 81.29 76.57 / 73.10 96.73 93.49 / 83.05 90.49 / 80.44
The results are taken from Vaze et al. (2022).
4.3 Performance on the Semantic Shift Benchmarks

### v2

@@ 4.3 Proposed Temperature Schedules
Instead of a constant value, we propose to schedule the temperature (replacing $\tau$ by $\mathcal{T}(e)$) from a range $\tau^{-},\tau^{+}]$ as the training progresses using a generalized cosine schedule, which is defined as
[…]
@@
where value $k\pi$ represents the delay of cosine wave with respect to the starting epoch, $P$ is the period of the wave (Figure 2) and $k$ can be from $[0,1]$. $k=0$ reduces Eq. (5) to the regular cosine schedule (CosSch), proposed by Kukleva et al. (2023) for the self-supervised tasks on the long-tailed datasets. With CosSch, the model starts training with a higher temperature $\tau^{+}$ and goes to a lower temperature $\tau^{-}$.
$\mathcal{T}_{\text{CosSch}}(e;\tau^{+},\tau^{-},P)=\mathcal{T}_{\text{GCosSch}}(e;\tau^{+},\tau^{-},P,k=0)$ (6)
@@
Proposed Negative Cosine Schedule. We find that rather than using Eq. (6), it is beneficial for the task switching if we start with a lower temperature $\tau^{-}$ and move towards a higher value $\tau^{+}$. Starting with a lower temperature, the model provides priority to fewer neighbors, learning the coarse structure of representation space, resulting in a sharper decision boundary. The open set samples remain distributed and distant from any cluster of known classes (due to heavy penalization of slight dissimilarity for the lower $\tau^{-}$). As temperature increases, the model prioritizes more neighbors and gradually pulls the positive samples to its own cluster, refining on the coarse representation space. This makes within-class representations more compact and the decision boundary smoother while the core separation learned earlier is maintained. The open set samples are not pulled as tightly because their features are unknown to the model, maintaining the separation. This leads to a richer and potentially more generalized representation space for both open set and closed set performance.
The second half cycle (decreasing from a higher to a lower temperature) facilitates the exploration by refining the model again for the instance-specific features and a smooth transition for the restart of next periodic cycle. The periodic restart can help the model to find a better nearby solution to generalize more effectively. The model settles down better if the final few epochs maintain a higher temperature rather than follow the wave (epochs 500-600 in Figure 2). Figure 1 illustrates the concept with UMAP projection of representation spaces for different TSs.
@@
For $k=1$ in Eq. (5), the temperature starts with a lower value and goes to a higher one, looking like a negative cosine wave, hence the name negative cosine schedule (NegCosSch).
$\mathcal{T}_{\text{NegCosSch}}(e;\tau^{+},\tau^{-},P)=\mathcal{T}_{\text{GCosSch}}(e;\tau^{+},\tau^{-},P,k=1)$ (7)
@@
Our experiments demonstrate that NegCosSch surpasses both CosSch and GCosSch (with other values of $k\neq 1$) in OSR performance. Although initially aimed for OSR, NegCosSch also improves closed set classification, while being applicable to any model architecture and loss function, such as CE, SupCon and ARPL (Chen et al., 2021) and incurring no additional computational burden, as it only includes an epoch-dependent temperature in a loss function.
Choice of $P$ and Other Proposed Monotonic Schedules. We find that a single monotonic increase with the first half cycle of NegCosSch (termed as Monotonic-NegCosSch or M-NegCosSch), where $P=2E$, is a sufficiently good TS.
$\mathcal{T}_{\text{M-NegCosSch}}(e;\tau^{+},\tau^{-})=\tau^{-}+0.5(\tau^{+}-\tau^{-})(1-\text{cos}(e\pi/E));\forall e$ (8)
Even a linear or an exponential temperature increase performs better than the baseline constant temperature. The exact formulations of these appear in Appendix B. Otherwise, $P$ in Eq. (7) can be chosen by dividing $E$ by the number of cycles. We denote periodic NegCosSch as P-NegCosSch. From ablation studies, we observe that varying $P$ does not change the performance significantly.
Choice of $(\tau^{+},\tau^{-})$. For any good value of constant temperature $\tau$ (which can be chosen from hyperparameter tuning) in the SupCon loss, it is better to use NegCosSch by setting $\tau^{+}=\tau+\Delta,\tau^{-}=\tau-\Delta$ (or alternatively, $\tau^{+}=\tau+\Delta,\tau^{-}=\tau$) with the increment $\Delta\approx 0.1$ or $0.2$. For example, hyperparameter tuning on TinyImageNet provides us a high OSR performance for $\mathcal{T}_{\text{Const}}$ with $\tau=0.2$. Therefore, $\mathcal{T}_{\text{NegCosSch}}(\tau^{+}=0.4,\tau^{-}=0.1)$ or $\mathcal{T}_{\text{NegCosSch}}(\tau^{+}=0.3,\tau^{-}=0.2)$ are better choices than $\mathcal{T}_{\text{Const}}(\tau=0.2)$, $\mathcal{T}_{\text{CosSch}}(\tau^{+}=0.4,\tau^{-}=0.1)$ and $\mathcal{T}_{\text{CosSch}}(\tau^{+}=0.3,\tau^{-}=0.2)$. For the CE loss, we find $\tau^{+}=2\tau,\tau^{-}=\tau/2$ as a good choice because in the CE loss, the temperature scales the logits instead of similarities. Utilizing this, we reduce the search space of $(\tau^{+},\tau^{-})$.
@@
Inference. For CE loss, we use the model as is for inference. However for SupCon loss, we remove the projection layer and a linear classifier is trained for evaluation. We use the maximum logit based scoring rule for OSR score.
@@ 5 Results and Discussion
Benchmarks. Here, we present the performance with different TSs on the TinyImageNet and the SSBs. The SSBs are defined on three fine-grained datasets: the Caltech-UCSD-Birds (CUB) (Wah et al. (2011)), FGVC-Aircraft (Krause et al. (2013)) and Stanford Cars (SCars) (Maji et al. (2013)). For SSBs, the open set classes are divided into ‘Easy’ and ‘Hard’ splits by computing the semantic similarity (based on the labeled visual attributes) of each pair of classes, the details of which can be found in Vaze et al. (2022). The different difficulty levels along with more training classes make these datasets harder OSR benchmarks than the other ones. Most of the OSR research does not report results on the SSBs. Moreover, we report the performance of our NegCosSch on the CIFAR benchmarks– CIFAR10, CIFAR+10 and CIFAR+50– from literature with their details in the Appendix.
@@
Training Details. We mostly follow the experiment settings and design choices from Vaze et al. (2022). For TinyImageNet, we use a VGG32-like model and for SSBs, we use a ResNet50 model pretrained on the places365 dataset [^1] In spite of our efforts, we could not find the same pretrained model mentioned in Vaze et al. (2022) online. Therefore, we use the pretrained model from Zhou et al. (2017) trained on places365, which is completely unrelated to the SSBs.. We run each experiment with 5 random seeds and report the average results. We also include results on a vision transformer model in Appendix F.2.
We perform ablations on temperatures from $\mathbb{T}_{\text{SupCon}}=\{0.025,0.05,0.1,0.2,0.3,0.4,0.5\}$ for the SupCon loss and from $\mathbb{T}_{\text{CE}}=\{0.5,1.0,2.0\}$ for the CE loss. However following Section 4.3, for SSBs, we fix $(\tau^{+},\tau^{-})$ at $(0.3,0.1)$ and $(0.4,0.1)$ for TinyImageNet with TSs and $\tau=0.2$ for $\mathcal{T}_{\text{Const}}$ in the SupCon loss. For CE loss, with the base temperature being $1.0$, we set $(\tau^{+},\tau^{-})$ at $(0.5,2.0)$. Similarly, we perform an ablation on $P$ from $\{100,200,1200\}$; however, here we report results using $P=200$ for all periodic TSs and $P=2E=1200$ for M-CosSch and M-NegCosSch to ensure consistent comparison. Ablations on $(\tau^{+},\tau^{-})$, $P$ and the training details appear in the Appendix.
@@
Metrics. We report the closed set performance as a $C$-class classification using accuracy (%), the open set performance as known-unknown detection using AUROC (%), and the open set classification rate (OSCR %). OSCR measures the trade-off between the open set and the closed set performance (Dhamija et al. (2018)). Now, we discuss the results.
Table 1: Comparison of different TSs on CE loss. For SSBs, the OSR results are shown on ‘Easy/ Hard’ splits. We bold the top three results for each metric and underline the best case.
Accuracy (%) AUROC (%) OSCR (%) Accuracy (%) AUROC (%) OSCR (%)
Schedule CUB Aircraft
Constant (Baseline) 84.43 83.55 / 74.98 70.49 / 63.34 90.88 90.35 / 81.48 82.05 / 74.25
Linear decrease 81.64 79.86 / 71.75 65.15 / 58.59 90.58 89.53 / 79.7 81.08 / 72.41
Random 85.06 85.02 / 75.54 72.28 / 64.32 91 90.76 / 82.37 82.55 / 75.17
P-CosSch 84.63 84.5 / 74.24 71.51 / 62.93 90.8 90.04 / 81.81 81.76 / 74.51
M-CosSch 81.77 79.55 / 71.4 64.96 / 58.35 90.62 88.63 / 80.92 80.35 / 73.57
Logarithmic increase 85.15 84.91 / 76.07 72.25 / 64.82 91.19 90.86 / 82.58 82.77 / 75.47
Exponential increase (ours) 86.12 86.65 / 78.05 74.64 / 67.35 90.88 90.92 / 82.93 82.54 / 75.54
Linear increase (ours) 86.22 86.54 / 78.01 74.58 / 67.32 90.97 91.11 / 83.25 82.87 / 76
P-NegCosSch (ours) 86.3 86.85 / 77.6 74.89 / 67.01 91.33 91.41 / 83.15 83.43 / 76.14
M-NegCosSch (ours) 86.12 86.79 / 78.08 74.7 / 67.3 91.15 91.15 / 83.23 82.99 / 76
SCars TinyImageNet
Constant (Baseline) 96.76 94.03 / 84.82 91.04 / 82.19 84.55 82.85 74.74
Linear decrease 96.25 92.51 / 83.47 89.14 / 80.51 84.12 82.64 74.28
Random 97.06 94.31 / 85.27 91.58 / 82.85 83.51 78.37 69.88
P-CosSch 96.63 93.85 / 84.88 90.75 / 82.14 84.41 83.12 74.79
M-CosSch 96.27 92.21 / 82.7 88.84 / 79.73 84.19 82.76 74.36
Logarithmic increase 97.07 94.92 / 85.42 92.18 / 83.03 84.74 82.96 74.91
Exponential increase (ours) 97.27 95.08 / 86.03 92.5 / 83.75 84.98 83.05 75.15
Linear increase (ours) 97.19 95.19 / 86.18 92.55 / 83.86 84.9 83.16 75.19
P-NegCosSch (ours) 97.3 95.03 / 86.05 92.49 / 83.81 84.85 83.02 75
M-NegCosSch (ours) 97.22 95.18 / 86.26 92.57 / 83.95 84.24 82.79 74.41
5.1 Ablation Study on Temperature Schedules
Here, we compare the closed set and open set performance among different TSs, such as a random schedule, a linear decrease and an increase, exponential and logarithmic monotonic increases, periodic and monotonic CosSch and NegCosSch on CE loss. We present the results in Table 1. For most cases, our proposed TSs, such as P-NegCosSch, M-NegCosSch, linear and exponential increases perform better than the constant baseline, CosSchs and other listed schedules in terms of all metrics. In terms of a single best result across a column, P-NegCosSch wins at a maximum number (8 out of 18) of cases and M-NegCosSch and linear win at 4 cases each. The standard deviations of these results across the trials are presented in Table 3 in the Appendix.
Table 2: Performance on different OSR loss functions, with and without the proposed schedules. Open set results are shown on ‘Easy / Hard’ splits. We highlight the cases where our TS produces better results than the baseline and underline the best result for each case.
Accuracy (%) AUROC (%) OSCR (%) Accuracy (%) AUROC (%) OSCR (%)
Loss Schedule CUB Aircraft
CE (w/o LS) Constant 84.43 83.55 / 74.98 70.49 / 63.34 90.88 90.35 / 81.48 82.05 / 74.25
M-NegCosSch(ours) 86.12 86.79 / 78.08 74.7 / 67.3 91.15 91.15 / 83.23 82.99 / 76
P-NegCosSch(ours) 86.3 86.85 / 77.6 74.89 / 67.01 91.33 91.41 / 83.15 83.43 / 76.14
CE + LS (Vaze et al. (2022)) Constant 85.53 85.15 / 77.44 72.77 / 66.26 90.73 86.85 / 79.72 78.84 / 72.55
M-NegCosSch(ours) 86.21 87.66 / 79.06 75.53 / 68.23 91.34 88.25 / 81.19 80.62 / 74.36
P-NegCosSch(ours) 86.12 86.43 / 78.03 74.36 / 67.22 91.1 87.25 / 80.03 79.55 / 73.17
SupCon (w/o LS) Constant 83.43 86.94 / 73.95 72.42 / 61.66 90.71 88.78 / 81.79 80.51 / 74.37
M-NegCosSch(ours) 85.3 88.14 / 75.81 75.09 / 64.72 91.43 90.45 / 82.49 82.57 / 75.51
P-NegCosSch(ours) 84.12 87.5 / 74.95 73.54 / 63.13 90.61 89.27 / 81.97 80.96 / 74.55
SupCon + LS Constant 83.72 86.43 / 73.69 72.3 / 61.74 90.05 88.97 / 81.81 80.11 / 73.85
M-NegCosSch(ours) 85.28 88.05 / 75.78 74.97 / 64.63 90.55 89.47 / 81.85 80.95 / 74.28
P-NegCosSch(ours) 84.38 87.26 / 75.16 73.5 / 63.39 90.43 88.77 / 81.78 80.2 / 74.08
ARPL (Chen et al. (2021)) Constant 85.8 86.93 / 79.7 78.64 / 73.36 90.88 90.75 / 81.77 85.98 / 78.26
M-NegCosSch(ours) 86.47 87.6 / 80.53 79.65 / 74.41 91.26 91.55 / 82.01 86.65 / 78.39
P-NegCosSch(ours) 86.51 87.57 / 80.12 79.61 / 74.05 90.98 91.52 / 82.02 86.55 / 78.38
SCars TinyImageNet
CE (w/o LS) Constant 96.76 94.03 / 84.82 91.04 / 82.19 81.95 78.6 69.22
M-NegCosSch(ours) 97.22 95.18 / 86.26 92.57 / 83.95 81.98 79.21 69.84
P-NegCosSch(ours) 97.3 95.03 / 86.05 92.49 / 83.81 82.23 79.05 69.91
CE + LS Constant 97.05 94.67 / 84.35 91.95 / 82.02 84.55 82.85 74.74
M-NegCosSch(ours) 97.23 95 / 85.06 92.42 / 82.83 84.24 82.79 74.41
P-NegCosSch(ours) 97.23 94.82 / 84.54 92.24 / 82.31 84.85 83.02 75
SupCon (w/o LS) Constant 96.58 92.99 / 82.8 89.92 / 80.12 85.37 82.87 70.61
M-NegCosSch(ours) 96.79 93.57 / 82.76 90.66 / 80.26 85.4 83.21 70.98
P-NegCosSch(ours) 96.68 93.32 / 83.16 90.31 / 80.53 85.18 83.09 70.71
SupCon + LS Constant 96.6 93.03 / 83.32 89.95 / 80.63 85.18 82.65 70.31
M-NegCosSch(ours) 96.84 93.58 / 83.15 90.69 / 80.62 85.57 83.11 71.04
P-NegCosSch(ours) 96.69 93.45 / 83.29 90.43 / 80.66 85.23 83.05 70.72
ARPL Constant 97.37 95.22 / 85.89 93.46 / 84.7 85.02 83 74.89
M-NegCosSch(ours) 97.29 95.27 / 86.03 93.48 / 84.82 84.83 83.07 75.03
P-NegCosSch(ours) 97.21 95.25 / 85.71 93.47 / 84.52 85 83.12 75.07
5.2 Performance of our Temperature Schedules across Various Loss Functions